OpenStudy (crashonce):

2 litres of a liquid contains 60% alcohol. How much alcohol would be added to obtain a liquid containing 75% alcohol

OpenStudy (perl):

Can we can assume we are adding 100% alcohol? (.60 * 2 + 1.00x ) / ( 2 + x ) = .75

OpenStudy (perl):

the numerator is pure alcohol, and the denominator is the liquid

OpenStudy (crashonce):

dont quiteget it can u explain and not assume we are adding 100 alcohol

OpenStudy (perl):

it doesnt tell you what percentage alcohol you are adding, so i assumed it was 100% alcohol

OpenStudy (perl):

pure alcohol is 100% alcohol, im assuming

OpenStudy (perl):

unless there is a typo in the question

OpenStudy (danjs):

It says adding alcohol, that is alcohol 100%

OpenStudy (danjs):

keep track of units, that is more important

OpenStudy (perl):

i used the form pure alcohol / total solution = 75%

OpenStudy (danjs):

\[\frac{ 1.2 L Alcohol }{ 2.0 L Liquid }\]

OpenStudy (perl):

i divided each percent by 100 , to convert to decimal

OpenStudy (danjs):

yeah the equation is correct at the top, i'm just saying, the units are more important to keep track of, the numbers will fall out if you have your units end up correct.

OpenStudy (danjs):

as a general problem analysis

OpenStudy (perl):

ok ( .60 * 2 liters alcohol + 1.00x liters alcohol ) / ( 2 liters liquid + x liters of liquid ) = .75

OpenStudy (perl):

hard to fit in one line :)

OpenStudy (danjs):



OpenStudy (danjs):

Just the way i always did probs to make it easier

OpenStudy (perl):

\[\frac{ 1.2 L Alcohol + x }{ 2 L Alcohol + x }= 0.75\]

OpenStudy (crashonce):

@ganeshie8 do u have a simple solution ?

OpenStudy (perl):

In the numerator we have 1.2 liters of alcohol , plus the x liters of 100% alcohol that you add.

OpenStudy (radar):

Ima trying. If I let x = the amount of 75% solutionl....maybe it would be something like The water will be 25% of the solution. That means .8=.25x So the solution would be 3. 2 liters, and all additional solution would come from additional alcohol. I need to work on this more. I want to proof it to myself.

OpenStudy (radar):

Hey I think your formulae will work.

OpenStudy (radar):

75% of 3.2 would require 2.4 liters of alcohol, You already have 1.2 so you add 1.2 more. giving you 3.2 liter solution. In your formula what does x represents?

OpenStudy (radar):

It also gives a value of 1.2 for x, Yours is simpler and straightforward. I now follow thanks.

OpenStudy (radar):

The above was directed to @perl